OpenStudy (anonymous):

Solve the following equation analytically. 2^(x − 1) = 8 P

OpenStudy (anonymous):

Well, 2 to what power gives you 8?

OpenStudy (jdoe0001):

well, let's rewrite your equation like \(\bf 2^{x-1} = 8 \implies 2^{x-1} = 2^3\) so, what do you think about x-1?